OpenStudy (anonymous):

solve the equation. Separate by commas, if more than one answer. x^(2)9^(x)-2x9^(x)=3(9^(x))

OpenStudy (anonymous):

\[x ^{2}9^{x}-2x9^{x}=3(9^{x})\]

OpenStudy (anonymous):

Do this...

OpenStudy (anonymous):

Bring over the right-hand side and take out the common factor of 9^x. You have\[9^x(x^2-2x-3)=0\]

OpenStudy (anonymous):

That means

OpenStudy (anonymous):

either \[9^x=0\]or\[x^2-2x-3=0\]

OpenStudy (anonymous):

There is no x here that will make 9^x zero (if you plot any exponential function, you'll see that it never crosses below the x-axis)

OpenStudy (anonymous):

Thanks a lot for explaining the steps, much appreciated.

OpenStudy (anonymous):

But,\[x^2-2x-3=(x+1)(x-3)\]so your second factor has that\[(x+1)(x-3)=0 \rightarrow x=-1,3\]

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

feel free to become a fan ;)

OpenStudy (anonymous):

I will! :)

OpenStudy (anonymous):

how do I become a fan though?

OpenStudy (anonymous):

There should be a blue link next to my name...says, "Become a fan"

OpenStudy (anonymous):

nice, very nice